If we burn 4.00 g of methylhydrazine in a bomb calorimeter, the temperature of the calorimeter increases from 25.00ºC to 39.50 ° C. In a separate experiment, the heat capacity of the calorimeter was determined to be 7,794 kJ / ºC. Calculate the heat of reaction for combustion of one mole of CH6N2 in this calorimeter.


1
Expert's answer
2021-12-06T01:49:02-0500

39.5-25= 14.5

Molar mass of methylhydrazine=40.60

4/40.60= 0.0999

= 0.999×14.5×7794×4.184

= 42.237kJ